A cue ball with mass 170 g hits the stationary number 8 ball, which has mass 160 g. The cue ball and the 8 ball move with a speed of 0.2 m/s and 0.3 m/s respectively after the collision. What was the speed of the cue ball before the collision?
1
Expert's answer
2020-05-05T18:46:23-0400

applying momentum conservation

"P_i=P_f"

"170\times v=170\times0.2+160\times0.3=34+48=82"

v=0.48m/s